# Analysis and Synthesis Denoisers

### v2

@@ Abstract
In this work we study the behavior of the forward-backward (FB) algorithm when the proximity operator is replaced by a sub-iterative procedure to approximate a Gaussian denoiser, in a Plug-and-Play (PnP) fashion. In particular, we consider both analysis and synthesis Gaussian denoisers within a dictionary framework, obtained by unrolling dual-FB iterations or FB iterations, respectively. We analyze the associated minimization problems as well as the asymptotic behavior of the resulting FB-PnP iterations. In particular, we show that the synthesis Gaussian denoising problem can be viewed as a proximity operator. For each case, analysis and synthesis, we show that the FB-PnP algorithms solve the same problem whether we use only one or an infinite number of sub-iteration to solve the denoising problem at each iteration. To this aim, we show that each ”one sub-iteration” strategy within the FB-PnP can be interpreted as a primal-dual algorithm when a warm-restart strategy is used. We further present similar results when using a Moreau-Yosida smoothing of the global problem, for an arbitrary number of sub-iterations. Finally, we provide numerical simulations to illustrate our theoretical results. In particular we first consider a toy compressive sensing example, as well as an image restoration problem in a deep dictionary framework.
@@ 1 Introduction
find $\text{find }\widehat{x}\in\underset{x\in\mathbb{R}^{N}}{\operatorname{Argmin}}\;\frac{1}{2}\|Ax-y\|^{2}+p(x),$ (1.2)
@@
$\begin{array}[]{l}x_{0}\in\mathbb{R}^{N}\\ \text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}x_{k+1}=\operatorname{prox}_{\tau p}\big{(}x_{k}-\tau A^{*}(Ax_{k}-y)\big{)},\end{array}\right.\end{array}$ (1.3)
where $\operatorname{prox}_{\tau p}$ is the so-called proximity operator of $p$, and $\tau>0$ is a step-size chosen to ensure the convergence of the generated sequence ${(x_{k})}_{k\in\mathbb{N}}$ to a solution to (1.2). In (1.3), one has to compute the proximity operator of $p$ (defined in (1.8)) at each iteration. This operator can have an explicit formula for many simple choices of functions $p$ (see e.g. [16]). However this is not true in many cases of interest, and the proximity operator must be approximated numerically with sub-iterations. These sub-iterations can become computationally expensive in practice as they need to accurately approximate $\operatorname{prox}_{p}$. Recently, to improve the reconstruction quality and avoid these sub-iterations, proximal algorithms have also been paired with deep learning techniques for solving inverse imaging problems. Examples include deep dictionary learning (DDL) [50], Plug-and-Play (PnP) algorithms [7], and unfolded neural networks [25]. However, providing convergence guarantees for these methods is not straightforward.
A common prior’s choice is based on the sparsity of the unknown signal in a certain basis, with prior $p$ of the form of
@@
where $\lambda>0$ is a regularization parameter balancing data fidelity and regularization terms, $\Gamma\colon\mathbb{R}^{N}\to\mathbb{R}^{S}$ is a linear operator modeling a sparsifying transformed domain, and $g\colon\mathbb{R}^{S}\to\mathbb{R}^{N}$ is a convex, lower semi-continuous, proper function promoting sparsity in the transformed domain (e.g., $\ell^{1}$ norm). Notable choices for the sparsifying operator include the Fourier, DCT, or Wavelet transforms [37] as well as the Total Variation (TV) regularization [48, 8], depending on the properties of the target solution. The structure of the sought signal, through the operator $\Gamma$, can also be learned from a ground truth dataset in a supervised setting, leading to dictionary learning (DL) approaches [2, 36]. For such sparsity-based priors, there exists no close-form formula and one typically needs to resort to expansive sub-iterations. In this paper, we investigate the properties of such approaches, strategies to reduce the sub-iteration computational burden, and their convergence.
@@
We adopt a PnP framework, where the proximity operator of $p$ in (1.3) is replaced with a denoiser. Building our PnP algorithm with FB iterations, we investigate the behavior of the following FB-PnP scheme
$\begin{array}[]{l}x_{0}\in\mathbb{R}^{N}\\ \text{for }k=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}x_{k+1}=G\big{(}x_{k}-\tau A^{*}(Ax_{k}-y)\big{)},\end{array}\right.\end{array}$ (1.5)
@@
find $\text{find }\quad G_{\Gamma}(v)\approx\operatorname{prox}_{p}(v)=\underset{x\in\mathbb{R}^{N}}{\operatorname{argmin}}\;\frac{1}{2}\|x-v\|^{2}+g_{\lambda}(\Gamma x),$ (1.6)
@@
$\text{find}\quad G_{D}(v)\approx D\,\underset{z\in\mathbb{R}^{S}}{\operatorname{argmin}}\;\frac{1}{2}\|Dz-v\|^{2}+\lambda g(z),$ (1.7)
@@
Our first contribution is to analyze the minimization problems associated with the proposed AD and SD described in (1.6)-(1.7), when $\Gamma$ and $D$ are fixed dictionaries. In particular, we show that (1.7) corresponds to a proximity operator, and that the associated problem is equivalent to a synthesis formulation of problem (1.2)-(1.4). Second, we investigate the behavior of algorithm (1.5), where $G_{\Gamma}$ and $G_{D}$ are obtained with either $1$ sub-iteration of a FB-based scheme, or when a large enough number of these iterations are run for solving (1.6)-(1.7), leveraging a warm-restart strategy. We further investigate the behavior of a similar algorithm to (1.5) for solving a Moreau envelope smoothing of problem (1.2)-(1.4). In this context, we show that using a warm-restart strategy yield an algorithm which converges toward a solution of the smooth approximated problem, if the operators $G_{\Gamma}$ and $G_{D}$ satisfy some sufficient decrease conditions. Finally, we show through simulations that the proposed unrolling within PnP frameworks is well suited to adopt a DDL approach for learning dictionaries $\Gamma$ or $D$.
@@
The remainder of the paper is organized as follows. Section 2 first presents the denoiser based on the sparse coding of the analysis coefficient using a given dictionary. It relies on the so-called dual forward-backward algorithm. Then, the unrolled version of this denoiser is embedded in a PnP framework to solve general linear inverse problems for which we provide a convergence analysis. Section 3 gives a similar analysis for the synthesis-based approach. In particular, we show the equivalence between the global synthesis-based approach to solve a linear inverse problem and the PnP approach using a synthesis-based denoiser. Using a smooth, regularized version of the original problem, Section 4 provides an analysis of the PnP algorithm using analysis or synthesis-based denoiser exploiting a bi-level optimization framework. Finally, Section 5 illustrates the convergence properties of the two PnP approaches on a toy compressive sensing examples, as well as to an image restoration problem within a DDL framework.
@@
$(\forall v\in\mathbb{R}^{N})\quad\operatorname{prox}_{g}(v)=\operatorname*{arg\,min}_{x\in\mathbb{R}^{N}}g(x)+\frac{1}{2}\|x-v\|^{2}\ .$ (1.8)
Let $\mathcal{C}\subset\mathbb{R}^{N}$ be a closed, non-empty, convex set. We denote by $\iota_{\mathcal{C}}$ the indicator function of $C$, defined as, for every $x\in\mathbb{R}^{N}$, $\iota_{\mathcal{C}}(x)=0$ if $x\in\mathcal{C}$, and $\iota_{\mathcal{C}}(x)=+\infty$ otherwise. The Fenchel-Legendre conjugate function of $g$ is denoted by $g^{*}\in\Gamma_{0}(\mathbb{R}^{N})$ and is defined as, for every $v\in\mathbb{R}^{N}$, $g^{*}(v)=\sup_{x\in\mathbb{R}^{N}}\langle v,x\rangle-g(x)$. The Moreau’s identity is given by ${\operatorname{Id}\mkern-3.0mu=\mkern-3.0mu\operatorname{prox}_{g}\mkern-3.0mu+\mkern-3.0mu\operatorname{prox}_{g^{*}}}$, where $\operatorname{Id}$ denotes the identity operator. The infimal post-composition of $g$ by an operator $U\colon\mathbb{R}^{S}\to\mathbb{R}^{N}$ is defined as
An operator $G\colon\mathbb{R}^{N}\to\mathbb{R}^{N}$ is $\beta$-Lipschitz continuous with parameter $\beta>0$ if, for every $(x,y)\in{(\mathbb{R}^{N})}^{2}$, $\|G(x)-G(y)\|_{2}\leq\beta\|x-y\|$.
For further background on convex optimization, we refer the reader to [4, 46] and references therein.
2 Analysis denoiser for FB-PnP algorithm
In this section we focus on the analysis denoiser defined in (1.6), for a fixed dictionary $\Gamma$. By definition, the proximity operator of $g\circ\Gamma$ can be interpreted as a MAP estimate for a Gaussian denoising problem, where the least-squares function corresponds to the data-fidelity term, and $g\circ\Gamma$ is the penalization function. During the last decade, this interpretation has been leveraged to develop new hybrid optimization algorithms, dubbed PnP methods, where the proximity operator can be replaced by more powerful denoisers [7, 49] to solve inverse problems of the form of (1.1). For instance, the PnP formulation of the FB algorithm (1.3), called FB-PnP hereafter, is given by (1.5). More generally, PnP algorithms have been proposed in the literature using a wide range of proximal algorithms, including HQS algorithm, Douglas-Rachford algorithm, or primal-dual algorithms. Denoisers can be either hand-crafted (e.g. BM3D [18]), or learned (e.g. neural network denoisers [11, 47, 45]). Although PnP methods have shown outstanding performances in many applications, they can be unstable in practice. The main challenge is to guarantee that PnP iterations produce a converging sequence of estimates ${(x_{k})}_{k\in\mathbb{N}}$, without sacrificing reconstruction performances. In particular, it is well known that PnP algorithms output converging sequences ${(x_{k})}_{k\in\mathbb{N}}$ if the denoiser $G$ is firmly non expansive, as a consequence of fixed point theory [4]. Recently, they have been extensively studied, in particular when denoisers are neural networks [41, 52, 26, 27, 29, 43].
2.1 Analysis denoiser structure
As mentioned in Section 1, unless $\Gamma$ is (semi)-orthogonal, the AD problem (1.6) does not have an explicit solution. However, it can be solved by the FB algorithm when applied to the dual problem of (1.6) (in the sense of Fenchel-Rockafellar, see e.g. [4, Chap. 15]), as proposed in [14, 15]. The resulting dual-FB algorithm reads
$\begin{array}[]{l}\text{for }\ell=0,1,\ldots\\ \left\lfloor\begin{array}[]{l}u_{\ell+1}=\operatorname{prox}_{\sigma g_{\lambda}^{*}}\big{(}u_{\ell}-\sigma{\Gamma}(\Gamma^{*}u_{\ell}-v)\big{)}\end{array}\right.\end{array}$ (2.1)
where $u_{0}\in\mathbb{R}^{S}$ and $\sigma>0$ is a step-size. Then the following convergence result can be deduced from [14, Thm. 3.7].
@@ 2.1 Analysis denoiser structure
Assume that there exists $\underline{\sigma}>0$ such that ${\underline{\sigma}<\sigma<(2-\underline{\sigma})\|\Gamma\|_{S}^{-2}\ .}$ Then ${(u_{\ell})}_{\ell\in\mathbb{N}}$ converges to a solution $u^{\dagger}_{\Gamma}$ to the dual problem of (2.5), and $x^{\dagger}_{\Gamma}=v-\sigma\Gamma^{*}u^{\dagger}_{\Gamma}$ is a solution to (2.5).
algorithm (2.1) can be used as sub-iterations when included within the global FB algorithm (1.5), to approximate the computation of $\operatorname{prox}_{g_{\lambda}\circ\Gamma}$. However, for (1.5) to still benefit from convergence guarantees, the computation of the proximity operator must be very accurate, which can be difficult in practice as this is only achieved at asymptotic convergence of (2.1). In this section, we investigate the behavior of (1.5) with the AD, when only one iteration of algorithm (2.1) is computed.
@@
[…]
@@
$G^{\mathcal{A}}_{L,\lambda,v}(u_{0})=v-\Gamma\widetilde{G}^{\mathcal{A}}_{L,\lambda,v}(u_{0}).$ (2.4)
@@
    $(\forall u\in\mathbb{R}^{S})\quad T^{\mathcal{A}}_{\lambda,v}(u)=\operatorname{prox}_{\sigma g_{\lambda}^{*}}\left((\operatorname{Id}-\sigma{\Gamma}^{*}\Gamma)u+\sigma{\Gamma}^{*}v\right).$
    In [32], the authors describe the layers of Model 2.2 using a primal-dual formulation, where each layer is composed of two feed-forward layers, one acting in the dual domain, and one acting in the primal domain.
@@
    As a direct consequence of Theorem 2.1, it can be noticed that, for any $u_{0}\in\mathbb{R}^{S}$, if there exists $\underline{\sigma}>0$ such that $\underline{\sigma}<\sigma<(2-\underline{\sigma})\|\Gamma\|_{S}^{-2}$, then $\lim_{L\to+\infty}G^{\mathcal{A}}_{L,\lambda,v}(u_{0})=x_{\Gamma}^{\dagger}$, where $x_{\Gamma}^{\dagger}$ is the solution to the problem to
    find $\text{find }\quad x^{\dagger}_{\Gamma}=\operatorname{prox}_{g_{\lambda}\circ\Gamma}(v)=\underset{x\in\mathbb{R}^{N}}{\operatorname{argmin}}\;\frac{1}{2}\|x-v\|^{2}+g_{\lambda}(\Gamma x).$ (2.5)

### v3

@@ Abstract
In this work we study the behavior of the forward-backward (FB) algorithm when the proximity operator is replaced by a sub-iterative procedure to approximate a Gaussian denoiser, in a Plug-and-Play (PnP) fashion. Specifically, we consider both analysis and synthesis Gaussian denoisers within a dictionary framework, obtained by unrolling dual-FB iterations or FB iterations, respectively. We analyze the associated minimization problems as well as the asymptotic behavior of the resulting FB-PnP iterations. In particular, we show that the synthesis Gaussian denoising problem can be viewed as a proximity operator. For each case, analysis and synthesis, we show that the FB-PnP algorithms solve the same problem whether we use only one or an infinite number of sub-iteration to solve the denoising problem at each iteration. To this aim, we show that each ”one sub-iteration” strategy within the FB-PnP can be interpreted as a primal-dual algorithm when a warm-restart strategy is used. We further present similar results when using a Moreau-Yosida smoothing of the global problem, for an arbitrary number of sub-iterations. Finally, we provide numerical simulations to illustrate our theoretical results. In particular we first consider a toy compressive sensing example, as well as an image restoration problem in a deep dictionary framework.
@@ 1 Introduction
$\text{find }\widehat{x}\in\underset{x\in\mathbb{R}^{N}}{\operatorname{Argmin}}\;\frac{1}{2}\|Ax-y\|^{2}+p(x),$ (1.2)
@@
[…]
where $\operatorname{prox}_{\tau p}$ is the so-called proximity operator of $p$ (see (1.8) for its definition), and $A^{*}$ denotes the adjoint operator of $A$. In (1.3), $\tau>0$ is a step-size chosen to ensure the convergence of sequence ${(x_{k})}_{k\in\mathbb{N}}$ to a solution to (1.2). In (1.3), one has to compute the proximity operator of $p$ (defined in (1.8)) at each iteration. This operator can have an explicit formula for many simple choices of functions $p$ (see e.g. [19]). However this is not true in many cases of interest, and the proximity operator must be approximated numerically with sub-iterations. These sub-iterations can become computationally expensive in practice as they need to accurately approximate $\operatorname{prox}_{p}$. Recently, to improve the reconstruction quality and avoid these sub-iterations, proximal algorithms have also been paired with deep learning techniques for solving inverse imaging problems. Examples include deep dictionary learning (DDL), Plug-and-Play (PnP) algorithms, and unfolded neural networks (see, e.g., [55, 9, 32, 41, 28, 31] and references therein). However, providing convergence guarantees for these methods is not straightforward.
A common prior’s choice is based on the sparsity of the unknown signal in a certain basis. In this context, the prior $p$ takes the form of
@@
where $\lambda>0$ is a regularization parameter balancing data fidelity and regularization terms, $\Gamma\colon\mathbb{R}^{N}\to\mathbb{R}^{S}$ is a linear operator modeling a sparsifying transformed domain, and $g\colon\mathbb{R}^{S}\to\mathbb{R}$ is a convex, lower semi-continuous, proper function. The function $g$ is then chosen for promoting sparsity in the transformed domain (e.g., $\ell^{1}$ norm). Notable choices for the sparsifying operator $\Gamma$ include the Fourier, DCT, or Wavelet transforms [40] as well as the Total Variation (TV) regularization [53, 10], depending on the properties of the target solution. The structure of the sought signal, through the operator $\Gamma$, can also be learned from a ground truth dataset in a supervised setting, leading to dictionary learning (DL) approaches [2, 39]. For such sparsity-based priors, there exists no close-form formula and one typically needs to resort to expansive sub-iterations as well as inexact proximal methods allowing for computational errors to approximate the proximity operator (see, e.g., [20]). In this paper, we investigate the properties of such approaches, reducing the sub-iteration computational burden. We will in particular study convergence properties of the FB iterations (1.3) when the proximity operator is (very) roughly approximated.
@@
We adopt a PnP framework, where the proximity operator of $p$ in (1.3) is replaced by a denoiser. Building our PnP algorithm with FB iterations, we investigate the behavior of the following FB-PnP scheme
[…]
@@
$\text{find }\quad G_{\Gamma}(v)\approx\operatorname{prox}_{g_{\lambda}\circ\Gamma}(v)=\underset{x\in\mathbb{R}^{N}}{\operatorname{argmin}}\;\frac{1}{2}\|x-v\|^{2}+g_{\lambda}(\Gamma x),$ (1.6)
@@
$\text{find}\quad G_{D}(v)=Dz^{\dagger}_{D}\;\text{ where }\;z^{\dagger}_{D}\approx\underset{z\in\mathbb{R}^{S}}{\operatorname{argmin}}\;\frac{1}{2}\|Dz-v\|^{2}+g_{\lambda}(z),$ (1.7)
@@
Our first contribution is to analyze the minimization problems associated with the proposed AD and SD in (1.6)-(1.7) when $\Gamma$ and $D$ are fixed dictionaries. In particular we show that (1.7) corresponds to a proximity operator, and that the associated problem is equivalent to a synthesis formulation of problem (1.2)-(1.4) (see Section 3.1). Second, we investigate the behavior of algorithm (1.5), where $G_{\Gamma}$ and $G_{D}$ are obtained running $1$ sub-iteration of a FB-based scheme for solving (1.6)-(1.7), using a warm-restart strategy (see Sections 2 and 3, respectively). Specifically, we show that the iterations using the AD within a FB-PnP with only $1$ sub-iteration of a dual FB algorithm [17] are equivalent to those of the Loris-Verhoeven primal-dual algorithm [38]. Similarly, we show that the iterations using the SD within a FB-PnP with only $1$ sub-iteration of a FB algorithm are equivalent to the FB iterations for solving the full synthesis formulation of problem (1.2)-(1.4). We further investigate the behavior of a similar algorithm to (1.5) for solving a Moreau envelope smoothing of problem (1.2)-(1.4). In this context, we show that using a warm-restart strategy yields an algorithm which converges toward a solution of the smooth approximated problem, if the operators $G_{\Gamma}$ and $G_{D}$ satisfy some sufficient decrease conditions (see Section 4). Finally, we illustrate our theoretical study on two simulation examples: a toy compressive sensing example, and an example on image restoration. In particular, we show through simulations that the proposed framework is well suited to adopt a DDL approach for learning dictionaries $\Gamma$ or $D$ (see Section 5).
A summary of the methods investigated in this work is provided in Table 1, and their associated theoretical conditions for convergence are summarized in Table 2.
Inverse problem Variational problem Prox equivalence Operator / Algorithm Section
Analysis Denoiser (AD) $v=\overline{x}+\varepsilon_{d}w$ $\left\{\hskip-2.84544pt\begin{aligned} &\ x_{\Gamma}^{\dagger}=v-\Gamma^{*}u_{\Gamma}^{\dagger}\\ &\ u_{\Gamma}^{\dagger}=\underset{u\in\mathbb{R}^{S}}{\operatorname{argmin}}\;\frac{1}{2}\|\Gamma^{*}u-v\|^{2}+g^{*}_{\lambda}(u)\end{aligned}\right$. $x_{\Gamma}^{\dagger}=\operatorname{prox}_{g_{\lambda}\circ\Gamma}(v)$ $\widetilde{G}_{L,\lambda,v}^{\mathcal{A}}$ $=\underset{\text{$L$ compositions}}{\underbrace{T_{\lambda,v}^{\mathcal{A}}\circ\dots\circ T_{\lambda,v}^{\mathcal{A}}}}$ $T_{\lambda,v}^{\mathcal{A}}\colon u\in\mathbb{R}^{S}\mapsto$ $\operatorname{prox}_{\sigma g_{\lambda}^{*}}\big{(}u-\sigma{\Gamma}(\Gamma^{*}u-v)\big{)}$ 2.1
[…]
Synthesis Denoiser (SD) $v=\overline{x}+\varepsilon_{d}w$ $\left\{\hskip-2.84544pt\begin{aligned} &\ x_{D}^{\dagger}=Dz_{D}^{\dagger}\\ &\ z_{D}^{\dagger}=\underset{z\in\mathbb{R}^{S}}{\operatorname{argmin}}\;\frac{1}{2}\|Dz-v\|^{2}+g_{\lambda}(u)\end{aligned}\right$. $x_{D}^{\dagger}=\operatorname{prox}_{D\rhd g_{\lambda}}(v)$ $\widetilde{G}_{L,\lambda,v}^{\mathcal{S}}$ $=\underset{\text{$L$ compositions}}{\underbrace{T_{\lambda,v}^{\mathcal{S}}\circ\dots\circ T_{\lambda,v}^{\mathcal{S}}}}$ $T_{\lambda,v}^{\mathcal{S}}\colon u\in\mathbb{R}^{S}\mapsto$ $\operatorname{prox}_{\zeta g_{\lambda}}\big{(}z-\zeta{D^{*}}(Dz-v)\big{)}$ 3.1 & 3.2
[…]
Table 1: Summary of the different problems and associated algorithms investigated in this work. Notation are the same as those used throughout the article. Conditions for convergence of the FB-PnP iterations are summarized in Table 2.
Convergence conditions Convergence
Analysis denoiser (AD) within FB-PnP pour centrer $L\to\infty$ or $L=1$ $g$ convex, lsc, proper $0<{\tau}<2\|A\|_{S}^{-2}$ $0<{\sigma}<2\|\Gamma\|_{S}^{-2}$ verticalement Theorem 2.4 ($L\to\infty$) and Theorem 2.5 ($L=1$)
Synthesis denoiser (SD) within FB-PnP pour centrer $L\to\infty$ $g$ convex, lsc, proper $0<{\tau}<2\|A\|_{S}^{-2}$ $0<{\zeta}<2\|D\|_{S}^{-2}$ verticalement Theorem 3.7
Synthesis denoiser (SD) within FB-PnP pour centrer $L=1$ $g$ convex, lsc, proper $0<\tau\zeta<2\|AD\|_{S}^{-2}$ verticalement Theorem 3.8
Unfolded denoiser within FB-PnP pour centrer $L$ chosen such that (4.10) is satisfied $\tilde{g}$ convex, lsc, proper $\tilde{f}$ convex, Lipschitz-differentiable $\tau$ chosen according to Theorem 4.3(ii) verticalement Theorem 4.3
Table 2: Summary of the theoretical conditions necessary to ensure convergence of the algorithms investigated in this work, and summarized in Table 1. The symbol is used to highlight new results.
@@
The remainder of the paper is organized as follows. Section 2 is dedicated to the analysis-based approach. We first present the denoiser based on the sparse coding of the analysis coefficient using a given dictionary. Then, the unrolled version of this denoiser is embedded in a PnP framework to solve general linear inverse problems for which we provide a convergence analysis. Section 3 gives a similar study of the synthesis-based approach. In Section 4, we use a smooth regularized version of the original problem, to study the PnP algorithm with denoisers built with a finite number of iterations. Finally, Section 5 illustrates the convergence properties of the two PnP approaches on a toy compressive sensing example, as well as to an image restoration problem within a DDL framework.
@@
An element of $\mathbb{R}^{N}$ is denoted by $x={(x^{(n)})}_{1\leq n\leq N}$}. The standard Euclidean norm is denoted $\|\cdot\|$, the $\ell_{p}$ norm (for $p>0$) is denoted $\|\cdot\|_{p}$, and $\|\cdot\|_{S}$ denotes the spectral norm. We denote $\Gamma_{0}(\mathbb{R}^{N})$ the set of proper, lower semi-continuous convex functions from $\mathbb{R}^{N}$ to $(-\infty,+\infty]$. We use $\mathbb{N}^{*}:=\mathbb{N}\setminus\{0\}$ to denote the set of strictly positive integers, and $\mathbb{R}^{*}:=\mathbb{R}\setminus\{0\}$ for the set of nonzero real numbers. For a function $f\in\Gamma_{0}(\mathbb{R}^{N})$, we write $\min_{x\in\mathbb{R}^{N}}f(x)$ to denote the minimum value of $f$, and $\underset{x\in\mathbb{R}^{N}}{\operatorname{argmin}}\;f(x)$ for a unique minimizer, when it exists. When the minimizer is not unique, we use $\underset{x\in\mathbb{R}^{N}}{\operatorname{Argmin}}\;f(x)$ to denote the set of minimizers. Further, the notation $\underset{x\in\mathbb{R}^{N}}{\operatorname{minimize}}\;f(x)$ refers to the task of finding an element in $\underset{x\in\mathbb{R}^{N}}{\operatorname{Argmin}}\;f(x)$.
The proximity operator of a function $g\in\Gamma_{0}(\mathbb{R}^{N})$ is defined as [42]
$(\forall v\in\mathbb{R}^{N})\quad\operatorname{prox}_{g}(v)=\underset{x\in\mathbb{R}^{N}}{\operatorname{argmin}}\;g(x)+\frac{1}{2}\|x-v\|^{2}\ .$ (1.8)
Let $\mathcal{C}\subset\mathbb{R}^{N}$ be a closed, non-empty, convex set. We denote by $\iota_{\mathcal{C}}$ the indicator function of $C$ defined, for every $x\in\mathbb{R}^{N}$, as $\iota_{\mathcal{C}}(x)=0$ if $x\in\mathcal{C}$, and $\iota_{\mathcal{C}}(x)=+\infty$ otherwise. The Fenchel-Legendre conjugate function of $g$ is denoted by $g^{*}\in\Gamma_{0}(\mathbb{R}^{N})$ and is defined as, for every $v\in\mathbb{R}^{N}$, $g^{*}(v)=\sup_{x\in\mathbb{R}^{N}}\langle v,x\rangle-g(x)$. The Moreau’s identity is given by $\operatorname{Id}\mkern-3.0mu=\mkern-3.0mu\operatorname{prox}_{g}\mkern-3.0mu+\mkern-3.0mu\operatorname{prox}_{g^{*}}$}, where $\operatorname{Id}$ denotes the identity operator. The infimal post-composition of $g$ by an operator $U\colon\mathbb{R}^{S}\to\mathbb{R}^{N}$ is defined as
In particular, by definition of the $\inf$ operator, if, for $x\in\mathbb{R}^{N}$, the set $\{z\in\mathbb{R}^{S}|x=Uz\}$ is empty, then $U\rhd g=+\infty$. An operator $G\colon\mathbb{R}^{N}\to\mathbb{R}^{N}$ is $\beta$-Lipschitz continuous with parameter $\beta>0$ if, for every $(x,y)\in{(\mathbb{R}^{N})}^{2}$, $\|G(x)-G(y)\|\leq\beta\|x-y\|$.
For further background on convex optimization, we refer the reader to [4, 50, 51] and references therein.
2 Analysis denoiser for FB-PnP algorithm
In this section we focus on the analysis denoiser defined in (1.6), for a fixed dictionary $\Gamma$. By definition, the proximity operator of $g_{\lambda}\circ\Gamma$ can be interpreted as a MAP estimate for a Gaussian denoising problem, where the least-squares function corresponds to the data-fidelity term, and $g_{\lambda}\circ\Gamma$ is the penalization function. During the last decade, this interpretation has been leveraged to develop new hybrid optimization algorithms, dubbed PnP methods, where the proximity operator can be replaced by more powerful denoisers [9, 54] to solve inverse problems of the form of (1.1). For instance, the PnP formulation of the FB algorithm (1.3), called FB-PnP hereafter, is given by (1.5). More generally, PnP algorithms have been proposed in the literature using a wide range of proximal algorithms, including HQS algorithm, Douglas-Rachford algorithm, or primal-dual algorithms. Denoisers can be either hand-crafted (e.g. BM3D [21]), or learned (e.g. neural network denoisers [14, 52, 49]). Although PnP methods have shown outstanding performances in many applications, they can be unstable in practice. The main challenge is to guarantee that PnP iterations produce a converging sequence of estimates ${(x_{k})}_{k\in\mathbb{N}}$, without sacrificing reconstruction performances. In particular, it is well known that PnP algorithms output converging sequences ${(x_{k})}_{k\in\mathbb{N}}$ if the denoiser $G$ is firmly non expansive, as a consequence of fixed point theory [4]. Recently, they have been extensively studied, in particular when denoisers are neural networks [45, 57, 29, 30, 32, 47, 7].
We use the subscript $\Gamma$ in $x^{\dagger}_{\Gamma}$ and $u^{\dagger}_{\Gamma}$ to denote outputs obtained from denoising tasks using an analysis operator $\Gamma$, in the primal and dual domains respectively. Later, in Section 2.2, the subscript $\mathcal{A}$ will be used to refer to the solution of the full inverse problem using an AD built from $\Gamma$. This convention highlights the distinction between the denoising phase and the inverse problem phase.
2.1 Analysis denoiser structure
As mentioned in Section 1, unless $\Gamma$ is (semi)-orthogonal, the AD problem (1.6) does not have an explicit solution. However, it can be solved by the FB algorithm when applied to the dual problem of (1.6) (in the sense of Fenchel-Rockafellar, see e.g. [4, Chap. 15]), as proposed in [17, 18]. This problem aims to
$\text{find }u^{\dagger}_{\Gamma}{\in\underset{u\in\mathbb{R}^{S}}{\operatorname{Argmin}}\;}\frac{1}{2}\|\Gamma^{*}u-v\|^{2}+g_{\lambda}^{*}(u).$ (2.1)
The resulting dual-FB algorithm reads
@@ 2.1 Analysis denoiser structure
where $u_{0}\in\mathbb{R}^{S}$ and $\sigma>0$ is a step-size. Then the following convergence result can be deduced from [17, Thm. 3.7].
@@
Assume that $0<\sigma<2\|\Gamma\|_{S}^{-2}$. Then $(u_{\ell})$$_{\ell\in\mathbb{N}}$ converges to a solution $u^{\dagger}_{\Gamma}$ to the dual problem (2.1) of (1.6), and $x^{\dagger}_{\Gamma}=v-\Gamma^{*}u^{\dagger}_{\Gamma}$ is a solution to (1.6).
Algorithm (2.2) can be used as sub-iterations when included within the global FB algorithm (1.5), to approximate the computation of $\operatorname{prox}_{g_{\lambda}\circ\Gamma}$. However, for (1.5) to still benefit from convergence guarantees, the computation of the proximity operator must be accurate, which can be difficult in practice as this is only achieved at asymptotic convergence of (2.2). In this section, we investigate the behavior of (1.5) with the AD, when only one iteration of algorithm (2.2) is computed. In the remainder, we use the superscript $\mathcal{A}$ to identify operators associated with the analysis formulation.
@@
$\widetilde{G}^{\mathcal{A}}_{L,\lambda,v}(u_{0})=\underset{\text{$L$ compositions}}{\underbrace{T^{\mathcal{A}}_{\lambda,v}\circ\dots\circ T^{\mathcal{A}}_{\lambda,v}}}(u_{0}),$ (2.3)
@@
$G^{\mathcal{A}}_{L,\lambda,v}(u_{0})=v-{\Gamma^{*}}\widetilde{G}^{\mathcal{A}}_{L,\lambda,v}(u_{0}).$ (2.5)
@@
    $(\forall u\in\mathbb{R}^{S})\quad T^{\mathcal{A}}_{\lambda,v}(u)=\operatorname{prox}_{\sigma g_{\lambda}^{*}}\left((\operatorname{Id}-\sigma{\Gamma\Gamma^{*}})u+\sigma{\Gamma}v\right),$
    where $\operatorname{Id}-\sigma{\Gamma\Gamma^{*}}$ is a linear operator and $\sigma{\Gamma}v$ a bias. In [35], the authors describe the layers of Model 2.2 using a primal-dual formulation, where each layer is composed of two feed-forward layers, one acting in the dual domain, and one acting in the primal domain.
@@
    As a direct consequence of Theorem 2.1, it can be noticed that, for any $u_{0}\in\mathbb{R}^{S}$, if $0<\sigma<2\|\Gamma\|_{S}^{-2}$, then $\lim_{L\to+\infty}G^{\mathcal{A}}_{L,\lambda,v}(u_{0})=x_{\Gamma}^{\dagger}$, where $x_{\Gamma}^{\dagger}$ is the solution to Problem (1.6).